OpenStudy (anonymous):

Find all numbers x satisfying the given equation. abs(x-3) + abs(x-4) = 1

OpenStudy (asnaseer):

if you are looking for integer solutions then, as abs is always positive, you can only make 1 in these ways: 0 + 1 = 1 1 + 0 = 1 that should help you find suitable values for x

OpenStudy (anonymous):

how should i start this?

OpenStudy (asnaseer):

from the 1st equation I listed (i.e. 0 + 1 = 1) you get:\[|x-3|=0\]and\[|x-4|=1\]can you solve that 1st?

OpenStudy (foolaroundmath):

There are three intervals:- \( x > 4\), \(3 \le x \le 4\) and \(x < 3\) Note that \(|x| = x,\;\;\; x \ge 0\) \(= -x, \;x \le 0\) So, in the interval \(x > 4\), \(x-4 > 0\) and \(x -3 > 0\) i.e. \(x-3 + x-4 = 1 \implies 2x = 8\) or \(x = 4\) In the interval \(3 \le x \le 4\), \(x-4 < 0\) and \(x - 3 > 0\) i.e. \(x-3 + 4-x = 1 \implies 1 = 1\) which is always true. Thus all values of \(x \in [3,4]\) satisfy the equation. In the final interval \(x < 3\), \(x - 4 < 0\) and \(x - 3 < 0\) i.e. \(3-x + 4-x = 1 \implies 2x = 6\) or \(x=3\) So, the final solution is \(x \in [3,4]\) Any value between 3 and 4 (both inclusive) will work.